The Function g is defined below
g(x)=x^2-2x-3/x^2-81
Find all the values of x that are NOT in the domain of g
if there is more than one value, separate them with commas.

Answer by jim_thompson5910(35256)


Recall that dividing by zero is undefined, so it is not allowed.

Set the denominator equal to zero, factor and use the zero product property to find the x values that solve the equation

x^2 - 81 = 0
x^2 - 9^2 = 0
(x-9)(x+9) = 0
x-9 = 0 or x+9 = 0
x = 9 or x = -9

The two x values not in the domain are x = 9 and x = -9. These x values make the denominator x^2-81 equal to zero.